OpenStudy (anonymous):

Solve the equation for x by first making an appropriate substitution, such as u = e^x. 5e^2x-21e^x+4=0 x = (smaller value) x = (larger value)

OpenStudy (anonymous):

5u^2 - 21u + 4 = 0 (5u -1)(u - 4) = 0 u = 0.2 or u = 4

OpenStudy (anonymous):

e^x. = 4 then x = 1.386 e^x = 0.2 then x = -1.609

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

jimmyrep, you're awesome! thanks for all the help!

OpenStudy (anonymous):

no probs - tell my grand daughter that - shes on the picture with her cat

OpenStudy (anonymous):

haha. i wondering why ur name is jimmy but u have a pic of a female.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

wait how did u get the final answers? i cant seem to get them when i type them in my calc